# LEGO: Learning and Graph-Optimized Modular Tracker for Online Multi-Object Tracking with Point Clouds

### v4

@@ IV-B Implementation Details
During the training procedure, Adam with a momentum set as 0.09 and an initial learning rate of 0.1 is applied. The learning rate decay is 0.001, and the training is executed across 100 epochs. During the inference phase, the LiDAR scanning frequency is fixed at 10Hz, and the threshold for the existence probability provided by the object detector is set to 0. This means that all the objects provided by the object detector are kept as input to the tracker. Implementation details of the key modules are as follows:
@@
AMGN score calculation module: The number $K$ of AMGN module is set to 3.
@@ IV-C Performance Comparison with Other State-of-the-Arts
In this sub-section, the framework is evaluated and discussed with reference to a variety of MOT metrics.
@@ IV-C1 Quality of Detection Input
The performance of a tracker is inherently tied to the effectiveness of the integrated detector. In the conducted experiments, three distinct object detectors were assessed: the CasA 3D detector, the PointGNN 3D detector, and the PointRCNN 3D detector. These 3D detectors were chosen based on their performance in the KITTI 3D Object detection challenge and arranged in descending order of effectiveness.
IV-C2 Performance Comparison with State-of-the-Art Trackers Using LiDAR Only on 2D MOT Metrics
The typical tracking result of our LEGO has been visualized in Fig. 5. The first row and the second show the ground truth and tracking result reported by our LEGO in scene 0006, the track state for every object in the current frame is marked by a bounding box and its track ID. The third row and the fourth-row show results in scene 0008. As delineated in Table I, a series of comparative analyses were conducted on various tracking methodologies within the context of the KITTI tracking benchmark. First, we compare our method with PC3T [44], utilizing PointRCNN as the object detector. The results reveal that our method achieved an enhancement of 0.25 in HOTA and 0.37 in MOTA on the testing dataset. Further analysis is extended to other methodologies utilizing disparate detectors, as outlined in Table I. For instance, in a comparison with CenterTube [40], which employs the same detector PointGNN, our method demonstrate an improvement of 8.32 in HOTA and 1.17 in MOTA. Likewise, when compared with UG3DMOT [42] that utilizes the CasA detector, our method achieved an enhancement of 2.15 in HOTA and 2.63 in MOTA.
IV-C3 Performance Comparison with State-of-the-Art Trackers Using LiDAR and Camera on 2D MOT Metrics
Many tracking algorithms leverage the fusion of 2D camera images and 3D point clouds to optimize performance within the KITTI tracking benchmark. Our proposed method, LEGO, was rigorously evaluated against a variety of LiDAR and camera fusion-based trackers. The comparative analysis was performed in two key segments in Table II, based on the detectors employed. Firstly, LEGO was compared with trackers utilizing PointRCNN as their detector. In this context, a notable improvement was observed against StrongFusionMOT [57], with LEGO registering 2.4 increments in HOTA and 3.44 increments in MOTA. The second segment of the comparative analysis involved trackers that employ PointGNN as their detector. LEGO’s improvements were again manifest, with a 3.62 enhancement in HOTA over HIDMOT [59] and a 5.34 advancement in HOTA and 0.09 advancement in MOTA over DualTracker [55]. A specific comparison with IMSF-MOT [52] highlighted LEGO’s superior performance.
IV-C4 Comparison of Proposed Method and Other State-of-the-Art Trackers on 3D MOT Metrics
@@ IV-D Ablation Study for LEGO Modular Tracker
This section outlines the process of parameter-tuning for LEGO, the car class type is taken as an example.
IV-D1 Effectiveness of Weight between Matrix $\bm{A}$ and Matrix $\bm{B}$
Throughout the experimental phase, refinements were introduced to the cost weight $w_{B}$ to ascertain the optimal equilibrium between the geometry and motion cost matrix $\bm{A}$, and the AMGN score matrix $\bm{B}$. A comprehensive assessment was conducted across four distinct levels for $w_{B}$, specifically 0, 1, 2, 3, and 4, as described in Table IV. It was found that when $w_{2}$ was set to 2, the best outcomes were achieved. These results underscore the pivotal role played by the cost weight in determining the relative contributions of matrix $\bm{A}$ and matrix $\bm{B}$ to the association process.
TABLE IV: Ablation study on the weight $w_{B}$ for matrix $\bm{B}$ with proposed method on KITTI car validation dataset.
$w_{B}$ HOTA(%) AssA(%$\uparrow$ LocA(%$\uparrow$ MOTA(%$\uparrow$
@@ IV-D1 Effectiveness of Weight between Matrix $\bm{A}$ and Matrix $\bm{B}$
TABLE V: Ablation study on the threshold in track management with proposed method on KITTI car validation dataset.
$N^{t}$ HOTA(%) AssA(%$\uparrow$ LocA(%$\uparrow$
10 85.305 86.704 92.371
12 85.522 87.045 92.35
@@
15 84.411 88.452 92.469
24 84.189 88.198 92.457
34 83.692 87.466 92.337
IV-D2 Effectiveness of Threshold in Track Management
In the experiments, the $N^{t}$ parameter was adjusted as 14 to manage tracks. Table V provides the results obtained by varying the threshold. These findings indicate that an optimal threshold pre-processing operation, combined with the appropriate cost weight used in the proposed LEGO tracker, can improve the HOTA score.
@@ V Conclusion
In this paper, we propose an online LiDAR-based tracker, LEGO, that introduces the offset correction module and AMGN modules, and effectively integrates them into the existing tracking framework to tackle the inherent challenges of MOT. Our offset correction module demonstrates the capability to rectify certain errors in the detection results, contributing to improved tracking performance. Furthermore, the AMGN modules which use the learnable adjacency matrix to estimate the association relationships between detected objects and predicted objects in an efficient way. However, it is important to acknowledge that the use of graph structures in our approach may lead to longer training times due to increased computational demands. As a potential solution for future research, techniques, such as voxel downsampling could be implemented to reduce the parameter count in the GNN, thereby enhancing computational efficiency and reducing time consumption.

### v6

@@ IV-B Implementation Details
During the training, Adam with a momentum set as 0.09 and an initial learning rate of 0.1 is applied. The learning rate decay is 0.001, and the training is executed across 100 epochs. During the inference phase, the LiDAR scanning frequency is fixed at 10Hz, and the threshold for the existence probability provided by the object detector is set to 0. This means that all the objects provided by the object detector are kept as input to the tracker. The track management threshold $N_{t}$ and $M_{t}$ are set to 14 and 4, respectively. Implementation details of the key modules are as follows:
@@
AMGN score calculation module: The number of AMGN blocks, $K$, is set to 3.
Time Complexity: Our method achieves a frame rate of 27.83 FPS on a single NVIDIA RTX 3080 GPU with an Intel(R) Xeon(R) Platinum CPU, measured across the entire KITTI validation dataset. Although direct FPS comparisons with existing methods are challenging due to differences in hardware configurations, our tracking pipeline requires no post-processing operations, thus functioning fully online and in real-time. Compared to previously reported methods such as FANTrack [19] (25.0 FPS), 3DT [72] (33.3 FPS), mmMOT [20] (4.8 FPS), and GNN3DMOT [52] (5.2 FPS). Although the hardware configurations used by these methods were not disclosed, limiting direct comparisons, our method achieves 27.83 FPS, which meets the requirements for online tracking speed.
@@ IV-C Performance Comparison with Other State-of-the-Arts
In this subsection, the framework is evaluated and discussed concerning various MOT metrics.
@@ IV-C1 Quality of Detection Input
The performance of a tracker is inherently tied to the effectiveness of the integrated detector. Three distinct object detectors were assessed in the conducted experiments: the CasA 3D detector, the PointGNN 3D detector, and the PointRCNN 3D detector. These 3D detectors were chosen based on their performance in the KITTI 3D Object detection challenge and arranged in descending order of effectiveness.
IV-C2 Performance Comparison with State-of-the-Art Trackers Using LiDAR Only on KITTI dataset with 2D MOT Metrics
The typical tracking result of our LEGO has been visualized in Fig. 4. The first row and the second show the ground truth and tracking result reported by our LEGO in scene 0006, the track state for every object in the current frame is marked by a bounding box and its track ID. The third row and the fourth-row show results in scene 0008. As delineated in Table I, a series of comparative analyses were conducted on various tracking methodologies within the context of the KITTI tracking benchmark. First, we compare our method with PC3T [45], utilizing PointRCNN as the object detector. The results reveal that our method achieved an enhancement of 0.25 in HOTA and 0.37 in MOTA on the testing dataset. Further analysis is extended to other methodologies utilizing disparate detectors, as outlined in Table I. For instance, in a comparison with CenterTube [41], which employs the same detector PointGNN, our method demonstrates an improvement of 8.32 in HOTA and 1.17 in MOTA. Likewise, when compared with UG3DMOT [43] that utilizes the CasA detector, our method achieved an enhancement of 2.15 in HOTA and 2.63 in MOTA.
IV-C3 Performance Comparison with State-of-the-Art Trackers Using LiDAR and Camera on KITTI dataset with 2D MOT Metrics
Many tracking algorithms leverage the fusion of 2D camera images and 3D point clouds to optimize performance within the KITTI tracking benchmark. Our proposed method, LEGO, was rigorously evaluated against a variety of LiDAR and camera fusion-based trackers. The comparative analysis was performed in two key segments in Table II, based on the detectors employed. Firstly, LEGO was compared with trackers utilizing PointRCNN as their detector. In this context, a notable improvement was observed against StrongFusionMOT [62], with LEGO registering 2.4 increments in HOTA and 3.44 increments in MOTA. The second segment of the comparative analysis involved trackers that employ PointGNN as their detector. LEGO’s improvements are again manifest, with a 3.62 enhancement in HOTA over HIDMOT [64] and a 5.34 advancement in HOTA and 0.09 advancement in MOTA over DualTracker [60]. A specific comparison with IMSF-MOT [57] highlighted LEGO’s superior performance.
IV-C4 Comparison of Proposed Method and Other State-of-the-Art Trackers on KITTI dataset with 3D MOT Metrics
@@
IV-C5 Comparison of Proposed Method and Other State-of-the-Art Trackers on Waymo Dataset with 3D MOT Metrics
To demonstrate the generalization capability of our proposed LEGO tracker, the evaluation of LEGO tracker against other state-of-the-art methods on the Waymo dataset is also shown in Table IV. Using CenterPoint as the detector, LEGO demonstrates superior performance, achieving 58.30 MOTA for both L1 and L2 metrics. This represents a notable improvement of 1.4 over SimpleTrack on L2 MOTA. Furthermore, LEGO significantly outperforms previous methods, surpassing SimTrack and 3DMODT by 5.2 and 2.4 respectively on MOTA L1. LEGO also exhibits better performance with MOTP scores of 19.44 for both L1 and L2 metrics, demonstrating improvements of 2.04 and 0.54 compared to SimTrack and 3DMODT, respectively. When integrated with the CasA detector, LEGO’s performance improves with 63.1 MOTA and 23.6 MOTP.
IV-C6 Additional Analysis
Although our proposed LEGO tracker achieves superior HOTA and MOTA scores compared to other state-of-the-art methods (as shown in Table I), we observe an increased number of identity switches (IDS). The AMGN module primarily focuses on relational features derived from the current and immediately preceding frames. Without explicit long-term memory or robust temporal smoothing mechanisms, the tracker may face difficulties in maintaining consistent object identities over extended periods, especially within highly dynamic scenes, further increasing IDS. Moreover, the current absence of appearance-based features limits the stability of associations, exacerbating IDS issues. In future work, these limitations are planned to be addressed by incorporating a lightweight module or knowledge distillation integrated with appearance features, aiming to enhance the stability and robustness of associations and thereby reduce IDS.
@@ IV-D Ablation Study for LEGO Modular Tracker
This section outlines the ablation study for various key modules in LEGO, all the experiments are implemented on KITTI car validation dataset.
IV-D1 Effectiveness of Modules
[…]
IV-D2 Effectiveness of Weight between Matrix $\bm{A}$ and Matrix $\bm{B}$
Throughout the experimental phase, refinements were introduced to the cost weight $w_{B}$ to ascertain the optimal equilibrium between the geometry and motion cost matrix $\bm{A}$, and the AMGN score matrix $\bm{B}$. A comprehensive assessment was conducted across four distinct levels for $w_{B}$, specifically 0, 1, 2, 3, and 4, as described in Table VI. When $w_{B}$ is set to 0, the AMGN module is disabled, meaning the graph structure and associated learned relationships are not utilized. Consequently, this configuration yields the lowest tracking performance. It was found that when $w_{2}$ was set to 2, the best outcomes were achieved. These results underscore the pivotal role played by the cost weight in determining the relative contributions of matrix $\bm{A}$ and matrix $\bm{B}$ to the association process.
IV-D3 Effectiveness of Parameter $K$ in AMGN Module
[…]
TABLE V: Ablation study of proposed modules on the KITTI car validation dataset.
[…]
82.600 85.01 88.00
83.079 85.954 90.649
84.800 87.20 91.00
85.808 88.63 92.65
TABLE VI: Ablation study of the weight $w_{B}$ for matrix $\bm{B}$ with proposed method on KITTI car validation dataset.
[…]
@@ IV-D3 Effectiveness of Parameter $K$ in AMGN Module
TABLE VII: Ablation study on the parameter K in AMGN Module with proposed method on KITTI car validation dataset.
[…]
0 82.600 85.01 88.00
1 83.357 86.592 90.677
2 84.263 87.014 91.588
@@
4 85.189 88.038 92.257
@@ V Conclusion
In this paper, we propose an online LiDAR-based tracker, LEGO, that introduces the offset correction module and AMGN score calculation module, and effectively integrates them into the existing tracking framework to tackle the inherent challenges of MOT. Our offset correction module demonstrates the capability to rectify certain errors in the detection results, contributing to improved tracking performance. Furthermore, the AMGN blocks which use the learnable adjacency matrix to estimate the association relationships between detected objects and predicted objects in an efficient way. However, it is important to acknowledge that the use of graph structures in our approach may lead to longer training times due to increased computational demands. As a potential solution for future research, techniques, such as voxel downsampling could be implemented to reduce the parameter count in the GNN, thereby enhancing computational efficiency and reducing time consumption.